OpenStudy (anonymous):

solve using the quadratic formula: (x-3)(x+1)=-6

OpenStudy (anonymous):

first u have to multiply ad simplify thus we have x^2-2x-3=-6 or x^2-2x-3+6=0 or x^2-2x+3=0 x^2-2x+1 +3-1=0 (x-1)^2 +2=0 (x-1)^2-2i^2=0 (i=sqrt(-1)) (x-1+isqrt(2))(x-1-isqrt(2))=0 hence x=1+i sqrt(2) ,1-i sqrt(2)